# CReMa: Crisis Response through Computational Identification and Matching of Cross-Lingual Requests and Offers Shared on Social Media

### v1

@@ IV-D Analysis of a million-scale real-world dataset
TABLE XI: A sample of tweets in MegaGeoCOV classified as requests and offers.
@@
The MegaGeoCOV Extended dataset [29] comprised 33.9 million multi-lingual geo-tagged tweets. As the regular expressions and classifiers utilized for identifying requests and offers were tailored for English, non-English tweets were filtered out. Among the 17.8 million English tweets, 40% (7.2 million tweets) matched at least one regular expression, indicating potential requests and offers. These 7.2 million tweets underwent classification to identify exclusive requests and offers. Among them, our classifiers identified 210k tweets seeking assistance and 72k tweets offering help. A summary of the findings is presented in Table XII. Each classifier took less than 2 hours to classify 7.2 million tweets on an NVIDIA A100 GPU (80GB). Some request and offer tweets from the classification are listed in Table XI. We conducted a random sampling of 200 tweets categorized as requests and offers. Upon manual inspection, we observed that only 41% of the tweets classified as requests were actually soliciting help during the COVID-19 crisis. Similarly, among offer tweets, only 38% explicitly offered assistance. The remaining were more general requests and offers. Examples include:
@@
TABLE XII: Classification tasks on MegaGeoCOV Extended [29].
@@
We used the classification results to explore the distribution of crisis communications concerning requests and offer during the pandemic. The top contributors to the COVID-19 English discourse, included the United States, the United Kingdom, India, Canada, and Australia, collectively representing over 81% of the discussion. Among the top 10 countries, Australia had the highest ratio of offers to requests ($O:R$), at 0.53, indicating a potential difference in community engagement or resources available for assistance. Apart from the ones in Table XII, Canada had $O:R$ of 0.39, India 0.18, South Africa 0.33, Nigeria 0.19, Ireland 0.43, the Philippines 0.39, and Pakistan 0.30. We observed regional variations in $O:R$ within a country, attributed to factors such as population density and contemporaneous issues (refer Table XIII). In the US, Los Angeles had $O:R$ of 0.30, Manhattan 0.16, Chicago 0.33, Houston 0.4. For UK regions, South East had 0.09, Manchester 0.4, Edinburgh 0.34, London 0.3, Glasgow 0.55, and Birmingham 0.51. Similarly, for Australian cities, Melbourne 0.61, Sydney and Brisbane 0.57. We present global distribution of request and offer tweets for the countries with $>50$ tweets in Figure 3, and their respective $O:R$ in Figure 4. Results from $O:R$ suggest that maintaining limited $\delta_{distance}$ can be problematic for some regions as the number of requests are much higher compared to offers.
TABLE XIII: Top active regions in the US, the UK and Australia.
@@
IV-E Advancing the area
@@
V Conclusion
In this study, we proposed a systematic approach to integrate textual, temporal, and spatial features for the identification and matching of requests and offers shared on social media. We trained classifiers to identify if a text is asking for assistance, offering assistance, or irrelevant. Results showed that CrisisTransformers, a family of crisis-domain-specific pre-trained models, outperform strong baselines like RoBERTa, MPNet, and BERTweet in classification tasks of identifying request and offer texts. Texts classified as requests and offers are then processed through a sentence encoder to generate sentence embeddings. We used CrisisTransformers’ multi-lingual sentence encoder for this task. The encoder outperforms traditional embedding approaches such as word2vec and GloVe, as well as more sophisticated models such as Universal Sentence Encoder and Sentence Transformers. We show that a cross-lingual embedding space is effective in generating sentence embeddings required for the matching tasks where requests are matched with relevant offers. Further, we experimented with different vector search strategies and studied their effects on indexing/searching times and accuracy. Furthermore, we analyzed a million-scale geotagged tweets dataset to study crisis communications concerning requests and offers on social media during the COVID-19 pandemic.

### v2

@@ IV-D Analysis of a million-scale real-world dataset
TABLE XIII: A sample of tweets in MegaGeoCOV classified as requests and offers.
@@
The MegaGeoCOV Extended dataset [29] comprised 33.9 million multi-lingual geo-tagged tweets. As the regular expressions and classifiers utilized for identifying requests and offers were tailored for English, non-English tweets were filtered out. Among the 17.8 million English tweets, 40% (7.2 million tweets) matched at least one regular expression, indicating potential requests and offers. These 7.2 million tweets underwent classification to identify exclusive requests and offers. Among them, our classifiers identified 210k tweets seeking assistance and 72k tweets offering help. A summary of the findings is presented in Table XIV. Each classifier took less than 2 hours to classify 7.2 million tweets on an NVIDIA A100 GPU (80GB). Some request and offer tweets from the classification are listed in Table XIII. We conducted a random sampling of 200 tweets categorized as requests and offers. Upon manual inspection, we observed that only 41% of the tweets classified as requests were actually soliciting help during the COVID-19 crisis. Similarly, among offer tweets, only 38% explicitly offered assistance. The remaining were more general requests and offers. Examples include:
@@
TABLE XIV: Classification tasks on MegaGeoCOV Extended [29].
@@
We used the classification results to explore the distribution of crisis communications concerning requests and offer during the pandemic. The top contributors to the COVID-19 English discourse, included the United States, the United Kingdom, India, Canada, and Australia, collectively representing over 81% of the discussion. Among the top 10 countries, Australia had the highest ratio of offers to requests ($O:R$), at 0.53, indicating a potential difference in community engagement or resources available for assistance. Apart from the ones in Table XIV, Canada had $O:R$ of 0.39, India 0.18, South Africa 0.33, Nigeria 0.19, Ireland 0.43, the Philippines 0.39, and Pakistan 0.30. We observed regional variations in $O:R$ within a country, attributed to factors such as population density and contemporaneous issues (refer Table XV). In the US, Los Angeles had $O:R$ of 0.30, Manhattan 0.16, Chicago 0.33, Houston 0.4. For UK regions, South East had 0.09, Manchester 0.4, Edinburgh 0.34, London 0.3, Glasgow 0.55, and Birmingham 0.51. Similarly, for Australian cities, Melbourne 0.61, Sydney and Brisbane 0.57. We present global distribution of request and offer tweets for the countries with $>50$ tweets in Figure 3, and their respective $O:R$ in Figure 4. Results from $O:R$ suggest that maintaining limited $\delta_{distance}$ can be problematic for some regions as the number of requests are much higher compared to offers.
TABLE XV: Top active regions in the US, the UK and Australia.
@@
V Conclusion
In this study, we proposed a systematic approach to integrate textual, temporal, and spatial features for the identification and matching of requests and offers shared on social media. We trained classifiers to identify if a text is asking for assistance, offering assistance, or irrelevant. Results showed that CrisisTransformers, a family of crisis-domain-specific pre-trained models, outperform strong baselines like RoBERTa, MPNet, and BERTweet in classification tasks of identifying request and offer texts. Texts classified as requests and offers are then processed through a sentence encoder to generate sentence embeddings. We used CrisisTransformers’ multi-lingual sentence encoder for this task. The encoder outperforms traditional embedding approaches such as word2vec and GloVe, as well as more sophisticated models such as Universal Sentence Encoder and Sentence Transformers. We show that a cross-lingual embedding space is effective in generating sentence embeddings required for the matching tasks where requests are matched with relevant offers. Next, we experimented with different vector search strategies and studied their effects on indexing/searching times and accuracy. Furthermore, we analyzed a million-scale geotagged tweets dataset to study crisis communications concerning requests and offers on social media during the COVID-19 pandemic.
V-A Advancing the area